OpenStudy (anonymous):

Question:Rationalize the denominator of √18 over √3 Answers: 6 2√3 over 3 √6 3√3 over 2

OpenStudy (angela210793):

multiply up and down with a sqr3 and u'll get sqr18*sqr3/3=sqr(3*3*2)*sqr3/3=3sqr6/3=sqr6=√6